OpenStudy (anonymous):

Solve for x: 3|2x − 2| + 6 = 18 Please help, thank you :)

OpenStudy (ranga):

3|2x − 2| + 6 = 18 3|2x − 2| = 18 - 6 = 12 |2x − 2| = 12/3 = 4 Whenever you have |A| = B, solve for two cases: A = B or A = -B |2x - 2| = 4 implies: 2x - 2 = 4 or 2x - 2 = -4 solve each equation.

OpenStudy (anonymous):

I got X= 3 and X=-1

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

Thank you so much for you help! Can you check a recent problem I worked on?

OpenStudy (ranga):

sure.

OpenStudy (anonymous):

|2x + 12| = 18 And I got X=3 X=-3

OpenStudy (ranga):

x = 3 is correct but not x = -3.

OpenStudy (anonymous):

Ohh, can you show me where I went wrong please?

OpenStudy (ranga):

whenever you see an absolute operator in an equation such as |A| = B you can get rid of the absolute bars by splitting the equation into two parts: A = B or A = -B

OpenStudy (ranga):

|2x + 12| = 18 implies: 2x + 12 = +18 or 2x + 12 = -18 2x + 12 = +18 2x = 18 - 12 = 6 x = 6/2 = 3 2x + 12 = -18 2x = -18 - 12 = -30 x = -30/2 = -15

OpenStudy (anonymous):

Thank you!

OpenStudy (ranga):

You are welcome.